OpenStudy (anonymous):

(Linear Equations) Write the equation in slope-intercept form for a line that passes through the points (2, -1) and (-3, -11).

OpenStudy (anonymous):

slope= (-1-(-11))/(2-(-3))=10/5=2 y-(-1)=2*(x-2) y+1=2x-4 y=2x-5